I need help solving 0.6(x-50)=18-0.3(40-10x)
Answer by midwood_trail(310)

0.6(x-50)=18-0.3(40-10x)
The first thing you want to do is remove the parentheses using the distributive rule.
On the left side:
0.6 times x = 0.6x
0.6 times -50 = -30
===================
On the right side:
-0.3 times 40 = 18 -12
-0.3 times -10x = 3x
====================
We now have the following equation to play with:

0.6x - 30 = 18 -12 + 3x
We now combine like terms:
0.6x - 30 = 6 + 3x
0.6x - 3x = 30 + 6
-2.4x = 36
x = 36/-2.4
x = -15
I forgot to include 18 in my first reply.
Did you follow now?